OpenStudy (anonymous):

If an equilateral triangle is circumscribe about a circle of radius 10 sq 3 cm, determine the side of the triangle.?

ganeshie8 (ganeshie8):

draw it out

OpenStudy (anonymous):

can you draw it for me please

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

you knw OA = OB = \(10 \sqrt{3}\) you're supposed to find AB any guess how to approach this ? :)

OpenStudy (anonymous):

wait =.=

OpenStudy (wolf1728):

I'd say the triangle area is the place to start

OpenStudy (anonymous):

=.= I dont have any idea Im sorry

ganeshie8 (ganeshie8):

Yes, thats one way, but there is a very easy way... if u knw 30-60-90 triangle properties or lil trig

OpenStudy (anonymous):

I know the 30 60 90 rule :)

ganeshie8 (ganeshie8):

cool then its a piece of cake for u

OpenStudy (wolf1728):

If that is an equilateral triangle, then the altitude is sqroot(3) * (base/2)

OpenStudy (anonymous):

@ganeshie8 for some reason its not more details on the drawing please =.=

ganeshie8 (ganeshie8):

lets wait for wolf to finish, his method seems interesting :)

OpenStudy (wolf1728):

I erased it all. I was trying to figure it out by the area formula.

ganeshie8 (ganeshie8):

we can find it by area formula altitude = 10sqrt{3} + 10sqrt{3}/2 = 30sqrt(3)/2

OpenStudy (anonymous):

I'm sorry to say my head isn't working =.= pls draw it =.=

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

altitude AD = OA + OD = \(10 \sqrt{3} + \frac{10 \sqrt{3}}{2}\)

ganeshie8 (ganeshie8):

once u knw the altitude of equilateral triangle, u should be able to find its side eh ?

OpenStudy (wolf1728):

alt = (side/2) * sqrt(3)

ganeshie8 (ganeshie8):

So, \(\large 15\sqrt{3} = (side / 2) * \sqrt{3}\)

OpenStudy (wolf1728):

I hate to be the bearer of bad news but if we look at the OP once again: If an equilateral triangle is circumscribe about a circle of radius 10 sq 3 cm, determine the side of the triangle.? THhat would seem to mean that the triangle SURROUNDS the circle. (see attached)

ganeshie8 (ganeshie8):

I agree lol we have been working the wrong problem :|